12.2.12 Question Help * The MINITAB output shown below is the result of a regression analysis for 22 homes relating y Answer parts a-c. selling price (in dollars) to x the size of the house (in square Constant Size Coef 9183 71.461 SE Coof 10063 6.448 0.913 11.083 0.372 0.000 Using the given data, determine the test statistic t. 11.083 What is the P-value? P-value 0.000 (Round to three decimal places as needed ) Use the P-value and a significance level of 0.05 to form a conclusion Reject Ho and conclude that the variables are dependent Fail to reject Ho and conclude that the variables could be independent O (Round to two decimal places as needed) 1.15 PM n the edit fields and then click Check Answer 1.13 PM 3 MacBook Air
Answer #1

As given in the results,

  • The test statistic t is 11.083
  • P value is 0.000

At a significance value of 0.05, we see that p value < 0.05 i.e. 0.000 < 0.05, thus we reject the null hypothesis and conclude that the variables are dependent.

b)

For the population slope,

SE Coef = 6.448

Coef = 71.461

Confidence level = 95%

Confidence interval = Coef of slope \pm Margin of error

Margin of error = SE * critical value - t

For n=22, df is n-2 = 20

Using the t table at 95% confidence level, at df =20,

t critical value = 2.086

Thus, Margin of error = 6.448 * 2.086 = 13.451

Confidence interval = 71.461 \pm 13.451 = 58.01 , 84.91